The time period of a simple pendulum is 2 s. What will be its length on the Earth?
-2 What will be its length on the Moon if g =g /6? where g = 10 m s .
1
Expert's answer
2020-08-19T12:56:47-0400

On the Earth


"T=2\pi\sqrt{\frac{l}{g}}\to l=\frac{gT^2}{4\pi^2}=\frac{9.81\cdot2^2}{4\cdot3.14^2}\approx1(m)"


On the Moon


"T=2\pi\sqrt{\frac{l}{g}}\to l=\frac{gT^2}{4\pi^2}="


"=\frac{(9.81/6)\cdot2^2}{4\cdot3.14^2}\approx0.166(m)"